# Automated Federated Pipeline for Parameter-Efficient Fine-Tuning of Large Language Models

### v1

@@ 25                         $C_{i,\max}^{r}\leftarrow C_{i,\max}^{r}-f_{c}(\mathcal{W}^{m}_{i},r)$
When the batch sizes $\mathbf{B}$ is determined, for the $i$-th edge server, FedPipe can obtain its computing resource constraint only related to rank via $C_{i,\max}^{r}=C_{i,\max}/b_{i}$ based on Eqn. (2e). To pick up the ranks and weights from the sorted weights according to the importance (see Section 3.3), FedPipe first starts from the most important weight, and then tries the rank value sequentially in $\mathcal{Q}^{m}$ from high to low. If the computing capability corresponding to the highest ranked weight is below the maximum computing constraint $C_{i,\max}^{r}$, it is selected. If not, the second-highest-ranked weight is used to evaluate. The above steps repeat for the other weights until sets of weights and ranks reach the computing constraint $C_{i,\max}^{r}$. To the end, the rank adjustment is iterated for each weight, and the optimal LoRA adapters at each edge server are configured dynamically during every training round. We summarize the whole process steps in Algorithm 1.
3.5. Quantization for Adapters’ Training
@@
In our design, we adopt the NormalFloat (NF) quantization built based on Quantile Quantization (Dettmers et al., 2022), which offers significant benefits in model compression, computational complexity reduction, accuracy preservation, and hardware compatibility (Dettmers et al., 2023). NF data type ensures that each quantization bin has an equal number of values for the input tensor. It estimates the quantile through the empirical cumulative distribution function, enhancing quantization accuracy.
Since pre-trained model weights usually follow a zero-mean normal distribution (Dettmers et al., 2023), we first transform all weights to a zero-mean normal distribution within $[-1,1]$ range via standard deviation. After that, we estimate the $12^{k}+$ 1 quantiles to obtain a $k$-bit quantile quantization data type and normalize the value of this data type into the $[-1,1]$ range. Finally, we quantize an input weight tensor by normalizing it into $[-1,1]$ range through absolute maximum rescaling. Once the weight range and data type range match, we estimate the $2^{k}$ values $q_{i}$ of the data type as follows
(8) ${q_{i}}=\frac{1}{2}\left[{Q\left({\frac{i}{2^{k}+1}}\right)}+{Q\left({\frac{i}{2^{k}+1}}\right)}\right]$
@@
Figure 12. Adapter aggregation of FedPipe framework.
3.6. LoRA Adapters Aggregation and Deployment
Current FL model aggregation aggregates the whole models, but for LLM, transferring them to the central server brings a significant communication overhead (see Section 2.2). Therefore, FedPipe only allows each edge server to transfer rank decomposition matrices $\mathbf{B}$ and $\mathbf{A}$ (see Section 3.2) instead of the whole models in order to reduce the communication overhead. However, to merge all kinds of rank decomposition matrices from varying edge servers, we still face two challenges in aggregation and updating, respectively. On the one hand, the diverse rank decomposition matrices with different ranks from edge servers cannot be aggregated directly (e.g., different sizes of two edge servers’ $\mathbf{B}$s cannot be added). On the other hand, even if the local models are aggregated, the central server recovers $\mathbf{B}$ and $\mathbf{A}$ for the edge server possibly leading to deterioration rather than improvement in the training performance. To combat above challenges, as illustrated in Figure 12, the tailored heterogeneous model aggregation solution comprises the following three stages.
@@ 3.6. LoRA Adapters Aggregation and Deployment
(9) $\triangle\mathbf{W}_{i}^{m{\kern 1.0pt},t}=\mathbf{B}_{i}^{m{\kern 1.0pt},t}\cdot\mathbf{A}_{i}^{m{\kern 1.0pt},t}$
@@
(10) $\triangle\mathbf{W}^{m{\kern 1.0pt},t}=\sum_{i=1}^{N}\frac{\left|D_{i}\right|/b_{i}}{{\textstyle\sum_{i=1}^{N}}\left|D_{i}\right|/b_{i}}\triangle\mathbf{W}_{i}^{m{\kern 1.0pt},t}$
@@
c) Local Model Update. Instead of updating $\mathbf{B}_{i}^{m{\kern 1.0pt},t}$ and $\mathbf{A}_{i}^{m{\kern 1.0pt},t}$ directly, the $i$-th edge server updates $\triangle\mathbf{W}_{i}^{m{\kern 1.0pt},t}$ dispatched by the central server. Moreover, recalling Section 3.5, according to the memory constraint of the $i$-th edge server, the local model is quantized, and hence, $\triangle\mathbf{W}_{i}^{m{\kern 1.0pt},t}$ is deemed to quantize for merging it into the local model. The updating policy is followed by
(11) ${\mathbf{W}_{i}^{m{\kern 1.0pt},t+1}}=\mathbf{W}_{i}^{m{\kern 1.0pt},t}+quant(\triangle\mathbf{W}^{m{\kern 1.0pt},t}),$
(12) ${\mathbf{B}_{i}^{m{\kern 1.0pt},t+1}}=\mathbf{\overline{B}}_{i}^{m},{\mathbf{A}_{i}^{m{\kern 1.0pt},t+1}}=\mathbf{\overline{A}}_{i}^{m}$
@@
where $\mathbf{\overline{A}}_{i}^{m}$ as well as $\mathbf{\overline{B}}_{i}^{m}$ are the reinitialized rank decomposition matrices, and $quant(\cdot)$ means the quantization function.
4. Implementation and Experimental Setup
Hardware and hyper-parameters: We conduct experiments utilizing 16 NVIDIA GeForce RTX 3090-Ti GPU. To ensure fair comparisons, we retain the consistent settings for FedPipe and benchmarks across different tasks. For TC task on the 20NEWS dataset, mini-batch size, learning rate, and maximum sequence length are set to 4, 0.1, and 256, respectively. Similarly, for the NLG task on the E2E dataset with GPT-2 medium, we set mini-batch size, learning rate, and maximum sequence length to 8, 0.0002, and 512, respectively. Considering the limited 24 GB memory of the NVIDIA GeForce RTX 3090-Ti, we reduce the mini-batch size to 4 while keeping other hyperparameters unchanged when training the GPT-2 large model. The number of selected participating edge servers is set to 15 and the rank is set to 4 by default unless specified otherwise. All edge servers run in synchronized mode (Ho et al., 2013).
@@ 4. Implementation and Experimental Setup
Models: We deploy two well-known LLMs, Bert (Kenton and Toutanova, 2019) and GPT-2 (Radford et al., 2019). The Bert model has been widely used for various natural language understanding tasks, including text classification, questioning and answering, and named entity recognition. GPT-2 is famous for its ability to generate coherent and relevant content, making it well suited for text generation tasks, such as text completion and language translation. The Bert model has 12 layers and 110 million parameters, whereas the GPT-2 medium and large models have 24 layers and 355 million parameters and 36 layers and 774 million parameters, respectively.
@@
Datasets and tasks: We evaluate the training performance of FedPipe for text classification (TC) tasks on 20NEWS dataset (Lang, 1995) and natural language generation (NLG) tasks on E2E dataset (Novikova et al., 2017). The 20NEWS dataset comprises approximately 20000 newsgroup documents, while the E2E dataset consists of around 42000 training, 4600 validation and 4600 test examples from the restaurant domain. We adopt the 20NEWS dataset to train the Bert model for the TC task, which outputs labels selected from predefined collections (e.g., politics, sports, and entertainment). Similarly, we train the GPT-2 model on the E2E dataset for the NLG task, aiming to transform various forms of inputs into human-readable natural language.
@@
Benchmarks: We compare FedPipe with three classical benchmarks: (1) Vanilla Fine-Tuning (FT): FT fine-tunes full parameters of LLMs on each edge server, which is the default fine-tuning technique in most NLP literature (Kenton and Toutanova, 2019). (2) Low-Rank Adaptation (LoRA): LoRA introduces low-rank trainable modules  (Hu et al., 2022) in parallel to weight matrices for LLM fine tuning. Training only low-rank trainable modules while freezing the rest of model parameters substantially reduces the number of trainable parameters without incurring additional inference latency. (3) FedAdapter: FedAdapter (Cai et al., 2023b) is an edge-cutting FL framework tailored for efficient LLM fine-tuning. It identifies the optimal adapter configuration during the training process, catering to the demands of fine-tuning LLMs in a privacy-preserving manner while minimizing the training costs. To ensure fair comparisons, all benchmarks adopt the same model aggregation algorithm (FedAvg (McMahan et al., 2017)) and random client sampling scheme, consistent with the default settings in prior FL literature (Cai et al., 2023b; Lin et al., 2022; Zhang et al., 2024).
5. Evaluation
@@
5.1. Overall Performance Evaluation
@@
Table 2. GPT-2 medium (M) and large (L) with different adaptation methods on the E2E NLG Challenge.
Figure 13. The converged accuracy for Bert and GPT-2 models under heterogeneous and ideal settings.
Figure 14. The number of trainable parameters for Bert and GPT-2 models under heterogeneous setting, respectively.
Figure 15. The convergence rate for Bert and GPT-2 models under heterogeneous and ideal settings.
Figure 16. Ablation Evaluation on GPT2-L model.
Converged accuracy. Figure 13 shows the comparison between FedPipe and the three other benchmarks in converged accuracy on Bert and GPT-2 models under heterogeneous and ideal settings. It is clear that the proposed FedPipe framework outperforms the other benchmarks in both settings. In the homogeneous setting, the FedPipe framework reaches comparable accuracy on Bert model and outperforms the FT, FedAdapter, and LoRA benchmarks for PPL by 0.22, 0.31, and 0.06, respectively. This superiority is attributed to tailored important weights identification, which prioritizes the fine-tuning of key weights, thereby enhancing the efficiency of fine-tuning LLMs and achieving higher converged accuracy. Moreover, due to the lack of adaptive rank adjustment for computing heterogeneity, all benchmarks exhibit notably inferior converged accuracy under ideal than heterogeneous settings. In contrast, the FedPipe framework automatically adjusts rank size based on heterogeneous computing resources and only experiences a slight increase in PPL, nearly 0.01, under heterogeneous than ideal settings. Moreover, the performance gap between FedPipe and the other benchmarks is much larger than that in ideal settings, further demonstrating the superior adaptability of FedPipe to heterogeneous scenarios.
Trainable parameters. Figure 14 presents the number of trainable parameters for deploying FedPipe and other benchmarks on Bert and GPT-2 models under heterogeneous and ideal settings. In the ideal setting, we observe that FedPipe framework and LoRA with fixed rank size exhibit identical number of trainable parameters. This is due to FedPipe’s capability in fine-tuning all four trainable weights and selecting the maximum rank for each weight without resource constraints. FedPipe has the fewest number of trainable parameters under heterogeneous setting, which is nearly 25% lower than that under the ideal setting. Deploying FedPipe on Bert and GPT-2 models results in approximately 0.28M and 0.37M trainable parameters, whereas the trainable parameters for FT and FedAdapter framework is over 380 and 2.5 times larger than FedPipe on Bert model and about 960 and 8 times on GPT-2 models, respectively. The reduction in trainable parameters significantly diminishes the computing and communication overheads of fine-tuning LLMs, facilitating the deployment of LLMs. It is noteworthy that FedPipe does not require iterative searching for the optimal depth and width of the adapter layer, reducing the time overhead of the training process. In contrast, FedAdapter selects adapters from three different configuration groups, whereas FedPipe only adapts an initialized adapter configuration, and thus substantially saves storage resources compared to FedAdapter.
Convergence rate. Figure 15 compares the convergence rate for FedPipe and the other three benchmarks under heterogeneous and ideal settings. In the ideal and heterogeneous setting, FedPipe exhibits the fastest convergence speed, outpacing FT, FedAdapter, and LoRA by factors of 6, 2.7, and 2 to achieve 3.6 PPL as well as 24, 18 and 2 to achieve 4.0 PPL on GPT-2 model, respectively. Moreover, FedPipe demonstrates comparable accuracy to FT, and is only lower by 0.23% on Bert model and even outperforms FT by 0.22 in the metric of PPL on GPT-2 model under ideal setting. This primarily stems from our design of the important layer identification strategy, which prioritizes the fine-tuning of crucial layers to reduce the under-training rate, thus expediting model training. Conversely, in the heterogeneous setting, FedPipe surpasses other benchmarks in convergence rate and accuracy due to its adaptive adjustment of LoRA’s rank size to accommodate the heterogeneous computing resources of clients. Due to the absence of heterogeneous LoRA adapter configuration, the convergence speed and accuracy of FT, FedAdapter, and LoRa rapidly deteriorate in heterogeneous settings. For the GPT-2 models, we also conducted a comprehensive comparison of the precision across various metrics under homogeneous and heterogeneous settings, which is summarized in Table 2.

### v2

@@ III-D Heterogeneous LoRA Adapters Configuration
When the batch sizes $\mathbf{B}$ is determined, for the $i$-th edge server, FedPipe can obtain its computing resource constraint only related to rank via $C_{i,\max}^{r}=C_{i,\max}/b_{i}$ based on Eqn. (2e). To pick up the ranks and weights from the sorted weights according to the importance (see Section III-C), FedPipe first starts from the most important weight, and then tries the rank value sequentially in $\mathcal{Q}^{m}$ from high to low. If the computing capability corresponding to the highest ranked weight is below the maximum computing constraint $C_{i,\max}^{r}$, it is selected. If not, the second-highest-ranked weight is used to evaluate. The above steps repeat for the other weights until sets of weights and ranks reach the computing constraint $C_{i,\max}^{r}$. At the end, the rank adjustment is iterated for each weight, and the optimal LoRA adapters at each edge server are configured dynamically during every training round. We summarize the whole process steps in Algorithm 1.
@@ 25                         $C_{i,\max}^{r}\leftarrow C_{i,\max}^{r}-f_{c}(\mathcal{W}^{m}_{i},r)$
III-E Quantization for Adapters’ Training
@@
In our design, we adopt the NormalFloat (NF) quantization built based on Quantile Quantization [61], which offers significant benefits in model compression, computational complexity reduction, accuracy preservation, and hardware compatibility [62]. NF data type ensures that each quantization bin has an equal number of values for the input tensor. It estimates the quantile through the empirical cumulative distribution function, enhancing quantization accuracy.
Since pre-trained model weights usually follow a zero-mean normal distribution [62], we first transform all weights to a zero-mean normal distribution within $[-1,1]$ range via standard deviation. After that, we estimate the $12^{k}+$ 1 quantiles to obtain a $k$-bit quantile quantization data type and normalize the value of this data type into the $[-1,1]$ range. Finally, we quantize an input weight tensor by normalizing it into $[-1,1]$ range through absolute maximum rescaling. Once the weight range and data type range match, we estimate the $2^{k}$ values $q_{i}$ of the data type as follows
${q_{i}}=\frac{1}{2}\left[{Q\left({\frac{i}{2^{k}+1}}\right)}+{Q\left({\frac{i}{2^{k}+1}}\right)}\right]$ (8)
@@
III-F LoRA Adapters Aggregation and Deployment
Figure 12: Adapter aggregation of FedPipe framework.
Current FL model aggregation aggregates the whole models, but for LLM, transferring them to the central server brings a significant communication overhead (see Section II-B). Therefore, FedPipe only allows each edge server to transfer rank decomposition matrices $\mathbf{B}$ and $\mathbf{A}$ (see Section III-B) instead of the whole models in order to reduce the communication overhead. However, to merge all kinds of rank decomposition matrices from varying edge servers, we still face two challenges in aggregation and updating, respectively. On the one hand, the diverse rank decomposition matrices with different ranks from edge servers cannot be aggregated directly (e.g., different sizes of two edge servers’ $\mathbf{B}$s cannot be added). On the other hand, even if the local models are aggregated, the central server recovers $\mathbf{B}$ and $\mathbf{A}$ for the edge server possibly leading to deterioration rather than improvement in the training performance. To combat above challenges, as illustrated in Fig. 12, the tailored heterogeneous model aggregation solution comprises the following three stages.
@@ III-F LoRA Adapters Aggregation and Deployment
$\triangle\mathbf{W}_{i}^{m{\kern 1.0pt},t}=\mathbf{B}_{i}^{m{\kern 1.0pt},t}\cdot\mathbf{A}_{i}^{m{\kern 1.0pt},t}$ (9)
@@
$\triangle\mathbf{W}^{m{\kern 1.0pt},t}=\sum_{i=1}^{N}\frac{\left|D_{i}\right|/b_{i}}{{\textstyle\sum_{i=1}^{N}}\left|D_{i}\right|/b_{i}}\triangle\mathbf{W}_{i}^{m{\kern 1.0pt},t}$ (10)
@@
c) Local Model Update. Instead of updating $\mathbf{B}_{i}^{m{\kern 1.0pt},t}$ and $\mathbf{A}_{i}^{m{\kern 1.0pt},t}$ directly, the $i$-th edge server updates $\triangle\mathbf{W}_{i}^{m{\kern 1.0pt},t}$ based on $\triangle\mathbf{W}^{m{\kern 1.0pt},t}$ dispatched by the central server. Moreover, recalling Section III-E, according to the memory constraint of the $i$-th edge server, the local model is quantized, and hence, $\triangle\mathbf{W}^{m{\kern 1.0pt},t}$ should be quantized for merging it into the local model. The updating policy is followed by
${\mathbf{W}_{i}^{m{\kern 1.0pt},t+1}}=\mathbf{W}_{i}^{m{\kern 1.0pt},t}+quant(\triangle\mathbf{W}^{m{\kern 1.0pt},t}),$ (11)
${\mathbf{B}_{i}^{m{\kern 1.0pt},t+1}}=\mathbf{\overline{B}}_{i}^{m},{\mathbf{A}_{i}^{m{\kern 1.0pt},t+1}}=\mathbf{\overline{A}}_{i}^{m}$ (12)
@@
where $\mathbf{\overline{A}}_{i}^{m}$ and $\mathbf{\overline{B}}_{i}^{m}$ are the reinitialized rank decomposition matrices, and $quant(\cdot)$ means the quantization function.
IV Implementation and Experimental Setup
@@
Models: We deploy two well-known LLMs, LLaMA2 [4] and GPT-2 [2]. The LLaMA2 model has been widely used for various natural language understanding tasks, including text generation, summarization, and questioning and answering. The GPT-2 is famous for its ability to generate coherent and relevant content, making it well suited for text generation tasks, such as text completion and language translation. The LLaMA2-7b model has 32 layers and 3.52 billion parameters, whereas the GPT-2 medium and large models have 24 layers and 355 million parameters and 36 layers and 774 million parameters, respectively.
@@
Datasets and tasks: We evaluate the training performance of FedPipe for instruction-following (IF) tasks on Alpaca dataset [63] and natural language generation (NLG) tasks on E2E dataset [64]. The Alpaca dataset comprises approximately 52,000 samples which spans diverse domains such as philosophy, economics, sociology, and science, while the E2E dataset consists of around 42,000 training, 4,600 validation and 4,600 test examples from the restaurant domain. We fine-tune the LLaMA2-7B model on the Alpaca dataset for IF task, which outputs detailed and contextually relevant responses based on given instructions. Similarly, we train the GPT-2 model on the E2E dataset for NLG task, aiming to transform various forms of inputs into human-readable natural language.
@@
Benchmarks: We compare FedPipe with three classical benchmarks: (1) Vanilla Fine-Tuning (FT): FT fine-tunes full parameters of LLMs on each edge server, which is the default fine-tuning technique in most NLP literature [44]. (2) Low-Rank Adaptation (LoRA): LoRA introduces low-rank trainable modules  [37] in parallel to weight matrices for LLM fine tuning. Training only low-rank trainable modules while freezing the rest of model parameters substantially reduces the number of trainable parameters without incurring additional inference latency. (3) FedAdapter: FedAdapter [43] is an edge-cutting FL framework tailored for efficient LLM fine-tuning. It identifies the optimal adapter configuration during the training process, catering to the demands of fine-tuning LLMs in a privacy-preserving manner while minimizing the training costs. To ensure fair comparisons, all benchmarks adopt the same model aggregation algorithm (FedAvg [53]) and random client sampling scheme, consistent with the default settings in prior FL literature [43, 65].
Hardware and hyper-parameters: We conduct experiments utilizing 16 NVIDIA GeForce RTX 3090-Ti GPU with the maximum memory of 24 GB. For resource heterogeneity, we constrain the each GPU’s maximum memory available for model training by randomly selecting values from a uniform distribution between 12 GB and 24 GB. To ensure fair comparisons, we retain the consistent settings for FedPipe and benchmarks across different tasks. For IF task on the Alpaca dataset, mini-batch size, learning rate, and maximum sequence length are set to 2, 0.0001, and 384, respectively. Similarly, for the NLG task on the E2E dataset with GPT-2 medium, we set mini-batch size, learning rate, and maximum sequence length to 8, 0.0002, and 512, respectively. Considering the limited 24 GB memory of the NVIDIA GeForce RTX 3090-Ti, we reduce the mini-batch size to 4 while keeping other hyperparameters unchanged when training the GPT-2 large model. The number of selected participating edge servers is set to 15 and the rank is set to 4 by default unless specified otherwise. All edge servers run in synchronized mode [66].
V Evaluation
@@
V-A Overall Performance Evaluation
(a) LLaMA2
(b) GPT-2
Figure 13: The converged accuracy for LLaMA2-7b and GPT-2 models under heterogeneous and ideal settings.
(a) LLaMA2
(b) GPT-2
Figure 14: The number of trainable parameters for LLaMA2-7b and GPT-2 models under heterogeneous setting, respectively.
(a) LLaMA2 (ideal).
(b) LLaMA2 (heterogeneous).
(c) GPT-2 (ideal).
(d) GPT-2 (heterogeneous).
Figure 15: The convergence rate for LLaMA2-7b and GPT-2 models under heterogeneous and ideal settings.
Converged accuracy. Fig. 13 shows the comparison between FedPipe and the three other benchmarks in converged accuracy on LLaMA2 and GPT-2 models under heterogeneous and ideal settings. It is clear that the proposed FedPipe framework outperforms the other benchmarks in both settings. In the homogeneous setting, the FedPipe framework reaches comparable accuracy on LLaMA2 model and outperforms the FT, FedAdapter, and LoRA benchmarks for PPL by 0.22, 0.31, and 0.06, respectively. This superiority is attributed to tailored important weights identification, which prioritizes the fine-tuning of key weights, thereby enhancing the efficiency of fine-tuning LLMs and achieving higher converged accuracy. Moreover, due to the lack of adaptive rank adjustment for computing heterogeneity, all benchmarks exhibit notably inferior converged accuracy under ideal than heterogeneous settings. In contrast, the FedPipe framework automatically adjusts rank size based on heterogeneous computing resources and only experiences a slight increase in PPL, nearly 0.01, under heterogeneous than ideal settings. Moreover, the performance gap between FedPipe and the other benchmarks is much larger than that in ideal settings, further demonstrating the superior adaptability of FedPipe to heterogeneous scenarios.
Trainable parameters. Fig. 14 presents the number of trainable parameters for deploying FedPipe and other benchmarks on LLaMA2 and GPT-2 models under heterogeneous and ideal settings. In the ideal setting, we observe that FedPipe framework and LoRA with fixed rank size exhibit identical number of trainable parameters. This is due to FedPipe’s capability in fine-tuning all four trainable weights and selecting the maximum rank for each weight without resource constraints. FedPipe has the fewest number of trainable parameters under heterogeneous setting, which is nearly 25% lower than that under the ideal setting. Deploying FedPipe on LLaMA2 and GPT-2 models results in approximately 7.69M and 0.37M trainable parameters, whereas the trainable parameters for FT and FedAdapter framework is over 350 and 4 times larger than FedPipe on LLaMA2 model and about 960 and 8 times on GPT-2 models, respectively. The reduction in trainable parameters significantly diminishes the computing and communication overheads of fine-tuning LLMs, facilitating the deployment of LLMs. It is noteworthy that FedPipe does not require iterative searching for the optimal depth and width of the adapter layer, reducing the time overhead of the training process. In contrast, FedAdapter selects adapters from three different configuration groups, whereas FedPipe only adapts an initialized adapter configuration, and thus substantially saves storage resources compared to FedAdapter.
Convergence rate. Fig. 15 compares the convergence rate for FedPipe and the other three benchmarks under heterogeneous and ideal settings. In the ideal and heterogeneous setting, FedPipe exhibits the fastest convergence speed, outpacing FT, FedAdapter, and LoRA by factors of 6, 2.7, and 2 to achieve 3.6 PPL as well as 24, 18 and 2 to achieve 4.0 PPL on GPT-2 model, respectively. Moreover, FedPipe demonstrates comparable performance to FT, and is only lower by 0.23 on LLaMA2 model and even outperforms FT by 0.22 in the metric of PPL on GPT-2 model under ideal setting. This primarily stems from our design of the important layer identification strategy, which prioritizes the fine-tuning of crucial layers to reduce the under-training rate, thus expediting model training. Conversely, in the heterogeneous setting, FedPipe surpasses other benchmarks in convergence rate and accuracy due to its adaptive adjustment of LoRA’s rank size to accommodate the heterogeneous computing resources of clients. Due to the absence of heterogeneous LoRA adapter configuration, the convergence speed and accuracy of FT, FedAdapter, and LoRa rapidly deteriorate in heterogeneous settings. For the GPT-2 models, we also conducted a comprehensive comparison of the precision across various metrics under homogeneous and heterogeneous settings, which is summarized in Table II.
@@ V-A Overall Performance Evaluation
TABLE II: GPT-2 medium (M) and large (L) with different adaptation methods on the E2E NLG Challenge.
(a) Important weights identification
(b) Heterogeneous configuration
(c) Quantization
Figure 16: Ablation Evaluation on GPT2-L model.
V-B Ablation Evaluation
Important weights identification (IWI). Fig. 16a shows the impact of the selection of trainable weights on GPT2-L model under heterogeneous setting. Considering that the average number of trainable weights in our training process is approximately 3, we compare the IWI method with a random combination of three trainable weights for fair comparison. The proposed IWI method exhibits lower PPL than all three random combinations of weights, which demonstrates the effectiveness of our IWI method. This is because large singular values usually correspond to the primary directions of variation in model weights, reflecting where the density of principal information is high. Therefore, it is efficient to prioritize fine-tuning the trainable weights with large singular values. Additionally, training performance on other metrics is summarized in Table III.